# SpiralMap: A Python library of the Milky Way’s spiral arms

### v1

@@
Mapping the structure of the Galaxy has been an ongoing endeavour for several decades, thanks to which we have come to piece together the components it is made up of such as the discs/bulge/halo and so on (Bland-Hawthorn & Gerhard, 2016). Of particular interest is also the signature of non-axisymmetry in the disc, principally present in the form of the central Galactic bar, the warp, and various spiral like features fanning across a large portion of the disc. Over the years, various groups have surveyed the Galaxy across wavelengths (Radio to optical) and have deduced the rich variety of spiral structure present in the disc. While some of the papers in literature provide machine readable data to trace out the spiral arms in their model, often this can be a cumbersome exercise for a user simply interested in extracting the coordinates and/or overplotting the spiral arms on another plot of interest, such as while comparing the locations of the arms to features in the velocity field (Khanna et al., 2023; Poggio et al., 2024, etc.). We show an example science application in Figure 4.
@@
Drimmel_Ceph_2024 Model based on Cepheid variables (Drimmel et al., 2024).
@@
1 Available models
In the current version, we have included the 8 models described below and summarised in Table 1. The arms in each model (except for contour maps) are given names for convenience.
  • Taylor_Cordes_1992: This model is bassed on radio and optical observations of H II regions. We use the model parameters presented in their Table 1 (Taylor & Cordes, 1993). There are four arms in this model (Arm1, Arm2, Arm3, Arm4).
@@ 1 Available models
  • Hou_Han_2014: This model is built upon the polynomial-logarithmic formulation introduced by Hou & Han (2014) and is based on a combination of H II, giant molecular clouds, and methanol MASER catalogs. Their model is publicly available. There are six arms in this model (Norma, Scutum-Centaurus, Sagittarius-Carina, Perseus, Local, Outer).
@@
2 Example Usage & Scientific application
As an example of a science case, we reproduce figures from (Khanna et al., 2024, hereafter K24) where SpiralMap was used. In particular, K24 constructed a model for the stellar density distribution in Red Clump stars in the Milky Way, and then compared the residuals of their best-fit models with the locations of non-axisymmetric structures such as spiral arms. Figure 4 shows two of their residual plots overlaid with spiral models (Drimmel_NIR_2000, Drimmel_Ceph_2024, & Reid_2019) using SpiralMap.

### v2

# arXiv:2506.11383v2[astro-ph.IM] 01 Mar 2026

@@
Mapping the structure of the Galaxy has been an ongoing endeavour for several decades, thanks to which we have come to piece together the components it is made up of such as the discs/bulge/halo and so on (Bland-Hawthorn & Gerhard, 2016). Of particular interest is also the signature of non-axisymmetry in the disc, principally present in the form of the central Galactic bar, the warp, and various spiral like features fanning across a large portion of the disc. Over the years, various groups have surveyed the Galaxy across wavelengths (Radio to Optical) and have deduced the rich variety of spiral structure present in the disc. While some of the papers in literature provide machine readable data to trace out the spiral arms in their model, often this can be a cumbersome exercise for a user simply interested in extracting the coordinates and/or overplotting the spiral arms on another plot of interest, such as while comparing the locations of the arms to features in the velocity field (Khanna et al., 2023; Poggio et al., 2025, etc.). We show an example science application in Figure 4.
@@
Vallee_1995 Model based on Galactic magnetic field/Dust/stars (Bower et al., 1995).
@@
Drimmel_Ceph_2024 Model based on Cepheid variables (Drimmel et al., 2025).
@@
I Available models
In the current version, we have included the 9 models described below and summarised in Table 1. The arms in each model (except for contour maps) are given names for convenience.
  • Taylor_Cordes_1992: This model is based on radio and optical observations of H II regions. We use the model parameters presented in their Table 1 (Taylor & Cordes, 1993). There are four arms in this model (Arm1, Arm2, Arm3, Arm4).
  • Vallee_1995: This model is based on a combination of the Galactic magnetic field, HI, HII, CO, and Dust distribution. We use their best model as depicted in their figure 3 (Bower et al., 1995). There are four arms in this model (Sagittarius, Scutum, Norma, Perseus).
@@ I Available models
  • Hou_Han_2014: This model is built upon the polynomial-logarithmic formulation introduced by Hou & Han (2014) and is based on a combination of H II, giant molecular clouds, and methanol MASER catalogs. Their model is publicly available. There are six arms in this model (Arm1, Arm2, Arm3, Arm4, Arm5, Arm6). Additionally their local arm is included as a regular logarithmic spiral.
@@
II Example Usage & Scientific application
As an example of a science case, we reproduce figures from (Khanna et al., 2025, hereafter K24) where SpiralMap was used. In particular, K24 constructed a model for the stellar density distribution in Red Clump stars in the Milky Way, and then compared the residuals of their best-fit models with the locations of non-axisymmetric structures such as spiral arms. Figure 4 shows two of their residual plots overlaid with spiral models (Drimmel_NIR_2000, Drimmel_Ceph_2024, & Reid_2019) using SpiralMap.